Circular No. 8199

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003hx IN NGC 2076 Further to IAUC 8198, J. Burket, M. Papenkova, and W. Li report the

LOTOSS discovery of an apparent supernova on unfiltered KAIT images taken on Sept. 12.5 (mag 14.3) and 13.5 UT (mag 14.4). SN 2003hx is located at α = 5^{h}46^{m}46^{s}.97, δ = −16^{o}47′00′′.6 (equinox 2000.0), which is 5′′.2 west and 2′′.6 south of the nucleus of NGC 2076. A KAIT image taken on Mar. 7.2 showed nothing at this position (limiting mag ∼ 18.5).

NOVA SCUTI 2003 S. Siviero, P. M. Marrese, and U. Munari, Asiago Observatory, write

that the spectroscopic evolution of the possible nova reported on IAUC 8190 and 8191, as monitored with the Asiago 1.22-m and 1.82-m telescopes, allows them to classify the object as a normal nova. Spectra (with AFOSC in échelle mode; range 360–860 nm; resolving power 3800) obtained on Sept. 6.83 UT display a well-developed emission-line spectrum dominated by Fe II lines (multiplets 27, 28, 37, 38, 40, 42, 48, 49, 73, and 74, among others), and Balmer lines characterized by a rapid decrement. Most lines display complex P-Cyg profiles, those of the Na I D lines having an emission width of 900 km/s and a terminal velocity of 1150 km/s for the absorption, with superimposed marked interstellar components. No nebular lines have so far developed.

H. Yamaoka, Kyushu University, reports his precise astrometry of this nova, from CCD images taken by B. Monard (on Aug. 30.7 UT) and by D. Rodriguez (Aug. 31.0), with respect to the UCAC-2 reference frame (∼ 100 stars; mean residual ∼ 0′′.1): α = 18^{h}49^{m}37^{s}.600, δ = −9^{o}33′50′′.85 (equinox 2000.0). Yamaoka adds that nothing is visible within 2′′ of this position on the red 1984 Digitized Sky Survey image or the 1999 2MASS public images. A. Henden, U.S. Naval Observatory, Flagstaff, reports the following position end figures from his images of the nova: 37^{s}.602, 50′′.76.

W. Hasubick, Buchloe, Germany, reports that nothing was visible at the position of the nova on a photograph taken with a 280-mm f/2.8 lens on Aug. 23.9 UT (limiting mag 14).

POSSIBLE NOVA IN M33 A. Shporer, E. O. Ofek, and T. Mazeh, Wise Observatory, Tel Aviv

University report further V magnitudes of the new possible nova in M33 (cf. IAUC 8195): Aug. 28.95 UT, [20:; Sept. 3.05, 16.9 ± 0.1.

2003 September 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green